OpenStudy (anonymous):

4x < 2x + 1 <= 3x + 2

OpenStudy (anonymous):

<= ??

OpenStudy (anonymous):

\[\le\]

OpenStudy (anonymous):

4x < 2x +1 4x-2x < 1 2x<1 x< 1/2 [ 2x + 1 \le 3x +1\] \[2x-3x \le 1-1\] \[-x \le 0\] \[x \ge 0\] \[0\le x < 1/2\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

can you help with another one?